OpenStudy (anonymous):

Please help! Given a=60, B=42degrees, C=58degrees. Find the area of triangle A B C. The area= ?

OpenStudy (anonymous):

Well, in this case we have 2 angles, and the side between. At least I think so. is that correct?

OpenStudy (anonymous):

Yes, I believe so

OpenStudy (anonymous):

so A = 80 degrees

OpenStudy (anonymous):

I'm not sure if there is a quicker way, but you can use the law of sines to get the other two sides.

OpenStudy (anonymous):

a/sin A = b/sin B = c/sin C Do you know that one?

OpenStudy (anonymous):

Oh o.k.

OpenStudy (anonymous):

Once you have the sides then you can find the length of an altitude to any one, or use Heron's formula again.

OpenStudy (anonymous):

Making sense?

OpenStudy (anonymous):

You don't need the law of sines.

OpenStudy (anonymous):

Yes. But how would I find the sides to fill in the formula?

OpenStudy (anonymous):

You do, sorry. We have the formula sin(A)/a = sin(B)/b We know A, B and a Calculate b

OpenStudy (anonymous):

Then use a and angle C to find the altitude down to b Use b and h to get the area.

OpenStudy (anonymous):

Ohhhh thanks! :)

OpenStudy (anonymous):

So sin(80)/60=.0164. Now to find altitude I... take 60 and 58?

OpenStudy (anonymous):

@telliott99

OpenStudy (anonymous):

back

OpenStudy (anonymous):

sin 80 deg = 0.985 sin 42 deg = 0.669 sin 58 deg = 0.848 60/0.985 = b/0.669 = c/0.848 b = 40.75 c = 51.65

OpenStudy (anonymous):

h/a = sin C h = 50.9 if I haven't made a dumb mistake.

OpenStudy (anonymous):

:/ I can't seem to match up: a. 1400 b. 2250 c. 1040 d. 1010

OpenStudy (anonymous):

I've given you the steps. Somewhere there is a mistake in the arithmetic, or perhaps (unlikely) the problem is in error.

OpenStudy (anonymous):

O.k. I'll firgure it out... Thanks!

OpenStudy (anonymous):

yw